OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2.

OpenStudy (anonymous):

Well we know that to find the slope, it's (y2-y1)/(2-x1)=slope So if (10,5) is your first point and (x,9) is your second and you already know the slope, just plug and chug! :)

OpenStudy (anonymous):

Oops, that should be (y2-y1)/(x2-x1)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]